OpenStudy (anonymous):

Urgent Help: if f(x)=x^4+ax^2+b has a stationary point (1,4) find: a and b insert (1,4) in the equation 4=1^4+a(1)^2+b 4=1+a+b a+b=3 a=3-b how to find the value of a and b?

OpenStudy (aum):

You then take the derivative, equate it to zero, put x = 1 and find 'a'.

OpenStudy (anonymous):

the derivative of f(x)?

OpenStudy (aum):

This is a calculus question, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the stationary point is f'(x)=0

OpenStudy (aum):

stationary point means the derivative is zero at that point.

OpenStudy (anonymous):

i mean the gradient

OpenStudy (aum):

Yes, the slope is zero at that point. If f'(c) = 0, then c is a stationary point.

OpenStudy (anonymous):

y'= 4x^3+0*x^2+2x*a+0

OpenStudy (anonymous):

y'= 4x^3+2xa

OpenStudy (anonymous):

4(1)^3+2(1)*a = 0

OpenStudy (aum):

correct. solve for 'a'

OpenStudy (anonymous):

am i doing it right?

OpenStudy (anonymous):

great

OpenStudy (anonymous):

4 +2a = 0 2a=-4 a=-2

OpenStudy (aum):

yes. put it in the other equation and find b.

OpenStudy (anonymous):

a=3-b -2=3-b b=5

OpenStudy (anonymous):

thank you so much for your help :)

OpenStudy (anonymous):

i really appreciate it

OpenStudy (aum):

You are welcome. I plotted f(x) here after putting in the values for 'a' and 'b'. As you can see, point (1,4) is on the curve and the tangent at (1,4) will be parallel to the x-axis with a slope of zero: https://www.wolframalpha.com/input/?i=x^4-2x^2%2B5&lk=4&num=1

OpenStudy (anonymous):

one more question... why let x=1 to find a?

OpenStudy (anonymous):

because i used x=1 in both equation f(x) and f'(x)

OpenStudy (aum):

(1,4) is a point on the curve. So the point should satisfy f(x). That is, 4 = f(1) (1,4) is a stationary point which implies the slope of f(x) is zero at (1,4) So f'(1) = 0

OpenStudy (anonymous):

Thanks :)

OpenStudy (aum):

If f'(c) > 0, the slope is positive at x = c and the function is increasing at x = c If f'(c) < 0, the slope is negative at x = c and the function is decreasing at x = c If f'(c) = 0, the slope is zero. The function is neither increasing nor decreasing at x = c. That is why it is called a "stationary point".

OpenStudy (anonymous):

if x=2 is the stationary point then what will y be?

OpenStudy (anonymous):

@aum

OpenStudy (aum):

Did you find x = 2 is a stationary point or are you just asking for some random x value?

OpenStudy (aum):

Either way, if you know x, all you have to do to find y is to put it in the function and evaluate f(x) when x = 2.

OpenStudy (anonymous):

there is another question that states if x^3+ax^2+bx has stationary points at x=2 and x=-4/3 find a and b?

OpenStudy (anonymous):

f'(2)=0?

OpenStudy (anonymous):

f'(-4/3)=0

OpenStudy (aum):

Exactly. You will get 2 equations and you can solve for the 2 unknowns: a and b.

OpenStudy (anonymous):

sweet ill try solve it and correct me if im wrong please

OpenStudy (aum):

okay.

OpenStudy (anonymous):

i got a different answer to what is in the textbook

OpenStudy (aum):

a = -1 and b = -8

OpenStudy (anonymous):

your answers are correct

OpenStudy (anonymous):

how were you able to answer it?

OpenStudy (aum):

f(x) = x^3+ax^2+bx f'(x) = 3x^2 + 2ax + b f'(2) = 3(2)^2 + 2a(2) + b = 0 12 + 4a + b = 0 4a + b = -12 ----- (1) f'(-4/3) = 3(-4/3)^2 + 2a(-4/3) + b = 0 3(16/9) - 8a/3 + b = 0 16/3 - 8a/3 + b = 0 multiply by 3 to get rid of fractions: 16 - 8a + 3b = 0 -8a + 3b = -16 ----- (2) Multiply (1) by 2 and add it to (2): 8a + 2b = -24 -8a + 3b = -16 add 5b = -40; b = -8. Put this in (1): 4a - 8 = -12 4a = -4 a = -1 a = -1, b = -8

OpenStudy (anonymous):

gosh i did a stupid mistake i inserted the value of x in f(x) instead of f'(x)

OpenStudy (anonymous):

now it makes sense

OpenStudy (aum):

you got to be careful

OpenStudy (anonymous):

exactly i shouldnt rush

OpenStudy (anonymous):

i understand what is a local maximum and a local minimum but what do they mean by "a negative point of inflection"?

OpenStudy (aum):

y = x^3+ax^2+bx a = -1, b = -8 y = x^3 - x^2 - 8x If you plot it, you will notice the stationary points are indeed at x = 2 and x = -4/3 http://prntscr.com/45dtae

OpenStudy (aum):

"point of inflection" is when f'(x) = 0 but the point is neither a maximum not a minimum. Such points will have the second derivative to be zero also. "c" could be a point of inflection if f'(c) = 0 and f''(c) = 0. If c is negative, I suppose you can call it a negative point of inflection.

OpenStudy (anonymous):

thank you again .. your explanations have been really helpful

OpenStudy (aum):

You are welcome. I have not heard the term "negative point of inflection" and so you may want to consult your book. But the above description is valid for "point of inflection"